# Robust-Wide: Robust Watermarking against Instruction-driven Image Editing

### v1

@@ Abstract
Instruction-driven image editing allows users to quickly edit an image according to text instructions in a forward pass. Nevertheless, malicious users can easily exploit this technique to create fake images, which could cause a crisis of trust and harm the rights of the original image owners. Watermarking is a common solution to trace such malicious behavior. Unfortunately, instruction-driven image editing can significantly change the watermarked image at the semantic level, making it less robust and effective.
We propose Robust-Wide, the first robust watermarking methodology against instruction-driven image editing. Specifically, we adopt the widely-used encoder-decoder framework for watermark embedding and extraction. To achieve robustness against semantic distortions, we introduce a novel Partial Instruction-driven Denoising Sampling Guidance (PIDSG) module, which consists of a large variety of instruction injections and substantial modifications of images at different semantic levels. With PIDSG, the encoder tends to embed the watermark into more robust and semantic-aware areas, which remains in existence even after severe image editing. Experiments demonstrate that Robust-Wide can effectively extract the watermark from the edited image with a low bit error rate of nearly 2.6% for 64-bit watermark messages. Meanwhile, it only induces a neglectable influence on the visual quality and editability of the original images. Moreover, Robust-Wide holds general robustness against different sampling configurations and other image editing methods such as ControlNet-InstructPix2Pix, MagicBrush, Inpainting and DDIM Inversion.
@@ 1 Introduction
Despite the success of the instruction-driven image editing technique, these models could be misused by malicious users. First, attackers can exploit these models to modify normal images to create fake news, causing a crisis of trust in an individual or even a country. Typical examples include changing someone's face or expression, forging endorsements for commercial gain, or taking off the clothes to produce vulgar images. Second, attackers can migrate the style based on a certain painting or make local modifications while keeping the basic composition of the painting unchanged to create new works, which shall be confirmed as plagiarism, infringing the IP rights of the work's owner. By integrating the personalization technique (e.g., Textual Inversion [11]), they may compose some concepts learned from other images in the edited image, which will undoubtedly lead to wider infringement on the concept's owners.
@@
We perform extensive experiments to demonstrate the robustness of Robust-Wide during the instruction-driven image editing process. It achieves a low Bit Error Rate (BER) of nearly 2.6% for 64-bit watermark messages, while preserving the visual quality and editability of original images. Besides the robustness against semantic distortions, Robust-Wide acquires inherent robustness against pixel-level distortions such as JPEG and color shifting, which are unseen during training. It also holds general robustness against different sampling configurations and models.
@@ 2 Background
Figure 1: The overall training pipeline of Robust-Wide.
@@ 2.1 Diffusion Model
Inspired by the non-equilibrium statistical physics, Diffusion Model (DM) [17] destroys the structure in a data distribution through an iterative forward diffusion process, and learns a reverse diffusion process to restore data's structure. Denoising Diffusion Probabilistic Model (DDPM) [18] further improves the performance of DM by training on a weighted variational bound with the folllowing objective:
@@ 2.2 Instruction-driven Image Editing
(2) ControlNet [19] is a dedicated framework that amplifies the capability and controllability of pre-trained T2I diffusion models by integrating spatial conditioning controls. Therefore, we can regard original images as spatial conditioning controls and train ControlNet on the dataset of InstructPix2Pix to realize instruction-driven image editing, which is called ControlNet-InstructPix2Pix [19]. (3) Afterwards, HIVE [8] improves InstructPix2Pix by harnessing human feedback to tackle the misalignment between editing instructions and resulting edited images. (4) MagicBrush [9] introduces the first large-scale and manually annotated dataset for instruction-guided real image editing and fine-tunes InstructPix2Pix on the dataset for better performance. (5) MGIE [10] uses multimodal large language models to derive expressive instructions and provide explicit guidance to further improve the editing performance while maintaining the efficiency.
In this paper, we mainly target InstructPix2Pix, and assess the generalization of our method to ControlNet-InstructPix2Pix and MagicBrush, since these models are open-sourced and publicly available.
@@ 2.3 Robust Watermarking
Robust watermarking is widely used for IP protection and forensics. Traditional methods (e.g., DWT-DCT [12] and DWT-DCT-SVD [12]) embed a pre-defined watermark message into transformed domains to achieve simple robustness against different operations, e.g., affine transformation, scaling, etc. HiDDeN [13] is the first to leverage deep neural networks for watermark embedding and extraction. Importantly, it simulates JPEG compression into a differential module, inserted into the famous encoder-noise layer-decoder framework for robustness enhancement. RivaGAN [20] applies a customized attention-based mechanism to embed diverse data and includes a separate adversarial networks for optimizing robustness. Afterwards, MBRS [14] utilizes a mixture of real JPEG, simulated JPEG, and noise-free images to get superior robustness performance against JPEG compression, which is also robust to various digital distortions. In addition to digital robustness, numerous works (LFM [21], StegaStamp [15], RIHOOP [22], PIMoG [16]) try to acquire robustness against physical operations such as screen-shooting, print-shooting, etc. However, none of them can resist instruction-driven image editing, which belongs to digital distortions.
In this paper, we compare our method with MBRS [14], the state-of-the-art digital-robust watermarking method. Furthermore, we adopt DWT-DCT [12], DWT-DCT-SVD [12], and RivaGAN [20] as the baseline methods, which are suggested by Stable Diffusion's official webpage [23].
@@ 3.2 Watermark Embedding
3.3 Partial Instruction-driven Denoising Sampling Guidance (PIDSG)
Existing watermarking solutions mainly consider the robustness against pixel-level distortions. In contrast, instruction-driven image editing changes an image significantly at the semantic level, making these approaches ineffective. We note that instruction-driven image editing involves the injection of a large number of different semantic instructions and varying degrees of image modifications at different semantic levels, which can be utilized to guide the robust and semantic-aware watermark embedding and extraction process. Thus, our intuitive idea is to incorporate the editing process into the end-to-end training framework. However, one challenge is that during the denoising sampling process of InstructPix2Pix, gradients are not allowed to flow directly. Introducing this process into training would result in the inability of gradients to propagate from the watermark decoder through the sampling process back to the watermark encoder. In other words, this would lead to a discontinuity in the computational graph, rendering the entire process non-differentiable. While a straightforward approach might be to open gradients for the numerous denoising sampling steps, this would introduce a significant memory overhead. To address this, we design the Partial Instruction-driven Denoising Sampling Guidance (PIDSG) module, which selectively allows gradients to flow only in the last $k$ sampling steps. This design not only makes the entire method feasible but also enables the process to be differentiable and amenable to end-to-end optimization.
@@ 4 Experiments
Method Image Size Watermark BER (%) $\downarrow$ PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$
Length (bits) w/o. Editing w/. Editing
@@
Robust-Wide (w/o. PIDSG) 512x512 64 0.0000 50.1558 55.3710 0.9982
Robust-Wide (w/. PIDSG) 512x512 64 0.0000 2.6579 41.9142 0.9910
Robust-Wide (w/o. PIDSG) $\dagger$ 512x512 64 0.0000 51.0801 55.3715 0.9983
Robust-Wide (w/. PIDSG) $\dagger$ 512x512 64 0.0000 2.6062 41.4038 0.9922
Table 1: Quantitative results compared with other methods. $\dagger$ denotes the results calculated on real-world images.
@@
BER (%) w/o. Editing w/. Editing
Original Images 49.8403 48.8550
Watermarked Images 0.0000 2.6579
Table 2: The integrity of Robust-Wide.
Images for Editing CLIP-Inormal-$\uparrow$ CLIP-Tnormal-$\uparrow$
Original Images 0.8402 0.2183
Watermarked Images 0.8430 0.2148
Table 3: The influence of Robust-Wide on the image editability.
Datasets. To train the embedding and extracting networks, we adopt 20k image-instruction pairs from the dataset used in InstructPix2Pix. We also select 1.2k additional samples that do not overlap with the above training data for evaluation by default. Besides, we collect some real-world images from the Internet, which cover 6 types (i.e., person, animal, object, architecture, painting, and scenery) and each type has 5 images. Then, we use InstructPix2Pix to edit these images based on 6 instructions and generate 8 images per instruction to finally obtain 1.44k edited images for testing.
We choose four baselines for comparisons, i.e., DWT-DCT [27], DWT-DCT-SVD [27], RivaGAN [20], and MBRS [14]. We directly use their official code for implementation. Noteably, the released model of MBRS only supports 256$\times$256 images with 256 bits.
@@
4.1 Effectiveness Evaluation
Table 1 compares the effectiveness of Robust-Wide with the baseline methods with different image sizes and watermark lengths. For Robust-Wide, we consider the implementation without and with PIDSG. From this table, it is obvious that none of the baseline methods can resist instruction-driven image editing, with the BER of around 50%. Comparably, Robust-Wide is effective with a low BER of 2.6579% in 1.2k InstructPix2Pix samples and 2.6062% in 1.44k real-world samples. Importantly, the removal of PIDSG leads to a complete failure, revealing the importance of PIDSG. Table 2 shows that Robust-Wide will not extract watermarks from original images with or without editing, guaranteeing forensic integrity.
4.2 Fidelity Evaluation
Table 1 also shows the PSNR and SSIM of different methods. We observe that Robust-Wide can achieve the comparable visual quality with other baseline methods. Table 3 compares the values of CLIP-I and CLIP-T for original and watermarked images. The slight difference indicates that Robust-Wide induces little influence on the editability. Figure 2 shows some visual results using our Robust-Wide, which further confirms the fidelity. Specifically, the normalized residual images are computed as $N(|I_{wm}-I_{ori}|)$, where $N(x)=(x-min(x))/(max(x)-min(x))$. From these images, it is evident that the watermark is predominantly embedded along the contours of the main subjects (such as people or objects) and in the background (secondary elements like buildings). We posit that these areas may be robust regions related to conceptual content and therefore Robust-Wide tends to embed watermarks into robust concept-aware areas.
4.3 Robustness Evaluation
Pixel-level Distortions. We apply different pixel-level distortions in three ways: (I) pre-processing watermarked images before editing; (II) post-processing watermarked images; (III) post-processing edited images based on watermarked images. Table 4 shows the extraction error of Robust-Wide against different pixel-level distortion types in three scenarios. It is obvious that Robust-Wide demonstrates strong robustness against these distortions even if we do not involve them during training.
Additionally, we also test Robust-Wide's robustness against DiffWA[28], a new watermark removal attack. This attack utilizes an image-to-image conditional diffusion model to add noise on the watermarked image and then restores the image while removing the embedded watermark. Due to the unavailability of DiffWA's source code, we utilize the open-source SDEdit[29] from Diffusers to carry out the process of adding noise and subsequently restoring the watermarked images. We configure the sampling parameters as follows: strength=0.2, prompt=Null, with all other sampling parameters as their default values. As shown in the ``Noise+Denoise" row of Table 4, Robust-Wide is effective in all three situations with BER$<10\%$.
Distortion Type BER(%)normal-$\downarrow$
I II III
None 2.6579 0.0000 2.6579
JPEG 2.7256 0.0013 2.7934
Median Blur 2.4926 0.0000 2.9574
Gaussian Blur 2.7074 0.0013 2.8489
Gaussian Noise 3.0672 0.0716 6.0546
Sharpness 2.6150 0.0013 2.7773
Brightness 12.2907 4.0733 9.4796
Contrast 5.1336 0.9210 4.3346
Saturation 3.0455 0.0000 3.0373
Hue 2.7591 0.0000 3.2829
Noise+Denoise 8.6401 3.9071 9.3654
Table 4: Robustness of Robust-Wide against pixel-level distortions in three processing scenarios.
@@ I II III
Different Sampling Configurations. We assess the robustness of Robust-Wide against different diffusion sampling configurations. Figure 3 shows the corresponding results. We have the following observations. (1) Robust-Wide is generally effective for different samplers. (2) As the total number of inference steps increases, the edited images become more fine-grained while the BER slightly increases (within a range of 1%). We explain that more details lead to larger differences between the edited and original images. (3) A larger text guidance scale ($s_{T}$) and smaller image guidance scale ($s_{I}$) indicate more severe image editing. Robust-Wide achieves BER of below 5% in all cases except when $s_{I}$ is 1, validating its general robustness in different settings.
Figure 4: Comparisons of image editing results among InstructPix2Pix, ControlNet-InstructPix2Pix, and MagicBrush.
@@
Figure 5: Robustness against Inpainting and DDIM Inversion.
@@
Other Image Editing Methods. In addition to InstructPix2Pix, we further evaluate our robustness against its extension ControlNet-InstructPix2Pix and MagicBrush on the 1.2k-samples dataset. Experiments show that Robust-Wide can achieve an average BER of 0.96% on ControlNet-InstructPix2Pix and 9.34% on MagicBrush. Figure 4 shows some visual examples. For the instruction ``swap sunflowers with roses", MagicBrush makes the edited image more different from the original image, leading to a relatively higher but still acceptable BER.
@@
We also tested the robustness of Robust-Wide against some other popular non-instruction-driven image editing methods, such as Inpainting and DDIM Inversion[30] based on Stable Diffusion Models. We found that, even though Robust-Wide has never seen these image editing methods during training, it still demonstrates effective resistance as shown in Figure 5.
@@
Continual Editing. A user may utilize InstructPix2Pix to perform multiple edits on a single image. An image edited by one user could also be spread to another for further editing, and this process can repeat several rounds. Hence, we need to ensure our watermarks can resist continual editing. As shown in Figure 6, we observe that the watermark embedded into the original image can be accurately extracted even after 3 editing rounds, demonstrating the watermark's strong robustness against continual editing.

### v3

@@ Abstract
Instruction-driven image editing allows users to quickly edit an image according to text instructions in a forward pass. Nevertheless, malicious users can easily exploit this technique to create fake images, which could cause a crisis of trust and harm the rights of the original image owners. Watermarking is a common solution to trace such malicious behavior. Unfortunately, instruction-driven image editing can significantly change the watermarked image at the semantic level, making current state-of-the-art watermarking methods ineffective.
To remedy it, we propose Robust-Wide, the first robust watermarking methodology against instruction-driven image editing. Specifically, we follow the classic structure of deep robust watermarking, consisting of the encoder, noise layer, and decoder. To achieve robustness against semantic distortions, we introduce a novel Partial Instruction-driven Denoising Sampling Guidance (PIDSG) module, which consists of a large variety of instruction injections and substantial modifications of images at different semantic levels. With PIDSG, the encoder tends to embed the watermark into more robust and semantic-aware areas, which remains in existence even after severe image editing. Experiments demonstrate that Robust-Wide can effectively extract the watermark from the edited image with a low bit error rate of nearly 2.6% for 64-bit watermark messages. Meanwhile, it only induces a neglectable influence on the visual quality and editability of the original images. Moreover, Robust-Wide holds general robustness against different sampling configurations and other popular image editing methods such as ControlNet-InstructPix2Pix, MagicBrush, Inpainting, and DDIM Inversion. Codes and models are available at https://github.com/hurunyi/Robust-Wide.
Keywords:
Watermarking Image Editing
@@ 1 Introduction
Despite the success of the instruction-driven image editing technique, these models could be misused by malicious users. First, attackers can exploit these models to modify normal images to create fake news, causing a crisis of trust in an individual or even a country. Typical examples include changing someone’s face or expression, forging endorsements for commercial gain, or taking off the clothes to produce vulgar images. Second, attackers can migrate the style based on a certain painting or make local modifications while keeping the basic composition of the painting unchanged to create new works, which shall be confirmed as plagiarism, infringing the IP rights of the work’s owner. By integrating the personalization technique (e.g., Textual Inversion [7]), they may compose some concepts learned from other images in the edited image, which will undoubtedly lead to wider infringement on the concept’s owners.
@@
We perform extensive experiments to demonstrate the robustness of Robust-Wide during the instruction-driven image editing process. It achieves a low Bit Error Rate (BER) of nearly 2.6% for 64-bit watermark messages, while preserving the visual quality and editability of original images. Besides the robustness against semantic distortions, Robust-Wide acquires inherent robustness against pixel-level distortions such as JPEG and color shifting, which are unseen during training. It also holds general robustness against different sampling configurations and even different popular editing models such as ControlNet-InstructPix2Pix, MagicBrush, Inpainting [10] and DDIM Inversion [16].
In summary, our contributions are as follows:
  • We point out the potential threat caused by the misuse of instruction-driven image editing and find current state-of-the-art watermarking methods are ineffective in the emerging case. In other words, we unearth a novel robustness requirement for current image watermarking.
  • We propose Robust-Wide, the first robust watermarking method for instruction-driven image editing. We introduce a novel Partial Instruction-driven Denoising Sampling Guidance (PIDSG) module, which forces the watermark embedded in the semantic-level rather than pixel-level.
  • Experimental results demonstrate that Robust-Wide can resist instruction-driven image editing, conventional pixel-level distortions, and different sampling configurations. More importantly, the proposed method exhibits robustness against a variety of popular editing models.
@@ 2.1 Diffusion Model
Inspired by the non-equilibrium statistical physics, Diffusion Model (DM) [24] destroys the structure in a data distribution through an iterative forward diffusion process, and learns a reverse diffusion process to restore data’s structure. Denoising Diffusion Probabilistic Model (DDPM) [9] further improves the performance of DM by training on a weighted variational bound with the folllowing objective:
@@ 2.2 Instruction-driven Image Editing
(2) ControlNet [31] is a dedicated framework that amplifies the capability and controllability of pre-trained T2I diffusion models by integrating spatial conditioning controls. Therefore, we can regard original images as spatial conditioning controls and train ControlNet on the dataset of InstructPix2Pix to realize instruction-driven image editing, which is called ControlNet-InstructPix2Pix [31]. (3) Afterwards, HIVE [32] improves InstructPix2Pix by harnessing human feedback to tackle the misalignment between editing instructions and resulting edited images. (4) MagicBrush [29] introduces the first large-scale and manually annotated dataset for instruction-guided real image editing and fine-tunes InstructPix2Pix on the dataset for better performance. (5) MGIE [6] uses multimodal large language models to derive expressive instructions and provide explicit guidance to further improve the editing performance while maintaining the efficiency. Besides, there are also some popular text-driven image editing methods such as Inpainting [10] and DDIM Inversion [16]. Inpainting edits an image by masking some regions of it and then regenerates the image according to the given text. DDIM Inversion [16] performs the reversed DDIM sampling process conditioned on the original text caption to convert the image to its partially noised version. Afterwards, the edited image is acquired by denoising the noised image given the edited text caption.
In this paper, we mainly target InstructPix2Pix, and assess the generalization of our method to ControlNet-InstructPix2Pix, MagicBrush, Inpainting and DDIM Inversion based on Stable Diffusion, since these models are open-sourced and publicly available.
@@ 2.3 Robust Watermarking
Robust watermarking is widely used for IP protection and forensics. Traditional methods (e.g., DWT-DCT [19] and DWT-DCT-SVD [19]) embed a pre-defined watermark message into transformed domains to achieve simple robustness against different operations, e.g., affine transformation, scaling, etc. HiDDeN [33] is the first to leverage deep neural networks for watermark embedding and extraction. Importantly, it simulates JPEG compression into a differential module, inserted into the famous encoder-noise layer-decoder framework for robustness enhancement. RivaGAN [30] applies a customized attention-based mechanism to embed diverse data and includes a separate adversarial network for optimizing robustness. Afterwards, MBRS [12] utilizes a mixture of real JPEG, simulated JPEG, and noise-free images to get superior robustness performance against JPEG compression. Similarly, CIN [14] combines invertible and non-invertible mechanisms for robustness to various digital distortions. In addition to digital robustness, numerous works (LFM [27], StegaStamp [25], RIHOOP [11], PIMoG [5]) try to acquire robustness against physical operations such as screen-shooting, print-shooting, etc. Recently, SepMark [28] utilizes a unified framework for source tracing and Deepfake detection which also achieves robustness against face manipulations, a special form of image editing. However, none of them can resist instruction-driven image editing (see Table 1.)
In this paper, we compare our method with current state-of-the-art watermarking methods such as MBRS [12], CIN [14], PIMoG [5], and SepMark [28]. We also adopt DWT-DCT [19], DWT-DCT-SVD [19], and RivaGAN [30] as the baseline methods, which are suggested by Stable Diffusion’s official webpage [26].
Figure 1: The overall training pipeline of Robust-Wide.
@@ 3.2 Watermark Embedding
3.3 Partial Instruction-driven Denoising Sampling Guidance
Existing watermarking solutions mainly consider the robustness against pixel-level distortions. In contrast, instruction-driven image editing changes an image significantly at the semantic level, making these approaches ineffective. We note that instruction-driven image editing involves the injection of a large number of different semantic instructions and varying degrees of image modifications at different semantic levels, which can be utilized to guide the robust and semantic-aware watermark embedding and extraction process. Thus, our intuitive idea is to incorporate the editing process into the end-to-end training framework. However, one challenge is that during the denoising sampling process (e.g., in InstructPix2Pix), gradients are not allowed to flow directly. Introducing this process into training would result in the inability of gradients to propagate from the watermark decoder through the sampling process back to the watermark encoder. In other words, this would lead to a discontinuity in the computational graph, rendering the entire process non-differentiable. While a straightforward approach might be to open gradients for the numerous denoising sampling steps, this would introduce a significant memory overhead. To address this, we design the Partial Instruction-driven Denoising Sampling Guidance (PIDSG) module, which selectively allows gradients to flow only in the last $k$ sampling steps. This design not only makes the entire method feasible but also enables the process to be differentiable and amenable to end-to-end optimization.
@@ 4 Experiments
Datasets. To train the embedding and extracting networks, we adopt 20k image-instruction pairs from the dataset used in InstructPix2Pix. We also select 1.2k additional samples that do not overlap with the above training data for evaluation by default. Besides, we collect some real-world images from the Internet, which cover 6 types (i.e., person, animal, object, architecture, painting, and scenery) and each type has 5 images which can be found in the supplementary material. Then, we use InstructPix2Pix [2] by default to edit these images based on 6 instructions and generate 8 images per instruction to finally obtain 1.44k edited images for testing.
We choose seven baselines for comparisons, i.e., DWT-DCT [19], DWT-DCT-SVD [19], RivaGAN [30], MBRS [12], CIN [14], PIMoG [5] and SepMark [28]. We directly use their official code for implementation. Noteably, the released model of MBRS and SepMark only support 256$\times$256 images with 256 bits and 128 bits separately and we used the Tracer of SepMark for watermark extraction after image editing. Since the released model of CIN only supports 128$\times$128 images and it is robust against the resize operation, so we resized the watermarked image to 256$\times$256 for image editing and then resized it back to 128$\times$128 for watermark extraction. PIMoG does not provide model weights, we trained the model ourselves using the official code.
Table 1: Quantitative results compared with other methods.
[…]
Length (bits) w/o Editing w/ Editing
@@
CIN [14] 128x128 30 0.0000 44.9888 40.3678 0.9763
PIMoG [5] 256x256 64 0.0000 49.9635 35.3183 0.9212
SepMark [28] 256x256 128 0.0084 28.1460 36.4341 0.9194
Robust-Wide 512x512 64 0.0000 2.6579 41.9142 0.9910
Robust-Wide 512x512 256 0.0000 4.1867 39.1842 0.9844
4.1 Effectiveness Evaluation
Table 2: The importance of PIDSG. $\dagger$ denotes the results on real-world images.
[…]
Length (bits) w/o Editing w/ Editing
Robust-Wide (w/o PIDSG) 512x512 64 0.0000 50.1558 55.3710 0.9982
Robust-Wide (w/ PIDSG) 512x512 64 0.0000 2.6579 41.9142 0.9910
Robust-Wide (w/o PIDSG) $\dagger$ 512x512 64 0.0000 51.0801 55.3715 0.9983
Robust-Wide (w/ PIDSG) $\dagger$ 512x512 64 0.0000 2.6062 41.4038 0.9922
Table 3: The integrity of Robust-Wide and the influence on the image editability.
Metrics w/o Editing (BER (%) $\downarrow$ ) w/ Editing (BER (%) $\downarrow$ ) CLIP-I $\uparrow$ CLIP-T $\uparrow$
Original Images 49.8403 48.8550 0.8402 0.2183
Watermarked Images 0.0000 2.6579 0.8430 0.2148
Table 1 compares the effectiveness of Robust-Wide with the baseline methods with different image sizes and watermark lengths. For Robust-Wide, we consider the implementation without and with PIDSG. From this table, it is obvious that none of the baseline methods can resist instruction-driven image editing, with the BER of around 50%. Comparably, Robust-Wide is effective with a low BER of 2.6579%. Table 2 also shows the effectiveness of Robust-Wide on 1.44k real-world samples. Importantly, the removal of PIDSG leads to a complete failure, revealing the importance of PIDSG. Moreover, Table 3 shows that Robust-Wide will not extract watermarks from original images with or without editing, guaranteeing forensic integrity.
4.2 Fidelity Evaluation
@@
Table 1 also shows the PSNR and SSIM of different methods. We observe that Robust-Wide can achieve the comparable visual quality with other baseline methods. Table 3 compares the values of CLIP-I and CLIP-T for original and watermarked images. The slight difference indicates that Robust-Wide induces little influence on the editability. Figure 2 shows some visual results using our Robust-Wide, which further confirms the fidelity (more results are provided in the supplementary material). Specifically, the normalized residual images are computed as $N(|I_{wm}-I_{ori}|)$, where $N(x)=(x-min(x))/(max(x)-min(x))$. From these images, it is evident that the watermark is predominantly embedded along the contours of the main subjects (such as people or objects) and in the background (secondary elements like buildings). We posit that these areas may be robust regions related to conceptual content and therefore Robust-Wide tends to embed watermarks into robust concept-aware areas.
4.3 Robustness Evaluation
Pixel-level Distortions.
Table 4: Robustness of Robust-Wide against pixel-level distortions.
BER(%) None JPEG Median Blur Gaussian Blur Gaussian Noise Sharpness Brightness Contrast Saturation Hue Noise+Denoise
I 2.6579 2.7256 2.4926 2.7074 3.0672 2.6150 12.2907 5.1336 3.0455 2.7591 8.6401
II 0.0000 0.0013 0.0000 0.0013 0.0716 0.0013 0.4733 0.9210 0.0000 0.0000 3.9071
III 2.6579 2.7934 2.9574 2.8489 6.0546 2.7773 9.4796 4.3346 3.0373 3.2829 9.3654
We apply different pixel-level distortions in three ways: (I) pre-processing watermarked images before editing; (II) post-processing watermarked images; (III) post-processing edited images based on watermarked images. Table 4 shows the extraction error of Robust-Wide against different pixel-level distortion types in three scenarios. It is obvious that Robust-Wide demonstrates strong robustness against these distortions even if we do not involve them during training.
Additionally, we also test Robust-Wide’s robustness against DiffWA[13], a new watermark removal attack. This attack utilizes an image-to-image conditional diffusion model to add noise on the watermarked image and then restores the image while removing the embedded watermark. Due to the unavailability of DiffWA’s source code, we utilize the open-source SDEdit[15] from Diffusers to carry out the process of adding noise and subsequently restoring the watermarked images. We configure the sampling parameters as follows: strength=0.2, prompt=Null, with all other sampling parameters as their default values. As shown in the “Noise+Denoise" row of Table 4, Robust-Wide is effective in all three situations with BER$<10\%$.
Different Sampling Configurations.
@@ 4.3 Robustness Evaluation
We assess the robustness of Robust-Wide against different diffusion sampling configurations. Figure 3 shows the corresponding results. We have the following observations. (1) Robust-Wide is generally effective for different samplers. (2) As the total number of inference steps increases, the edited images become more fine-grained while the BER slightly increases (within a range of 1%). We explain that more details lead to larger differences between the edited and original images. (3) A larger text guidance scale ($s_{T}$) and smaller image guidance scale ($s_{I}$) indicate more severe image editing. Robust-Wide achieves BER of below 5% in all cases except when $s_{I}$ is 1, validating its general robustness in different settings.
@@
Figure 4: General robustness against other editing methods such as InstructPix2Pix, ControlNet-InstructPix2Pix, MagicBrush, Inpainting, and DDIM Inversion.
@@
Other Popular Image Editing Methods. In addition to InstructPix2Pix, we further evaluate our robustness against its extension ControlNet-InstructPix2Pix and MagicBrush on the 1.2k-samples dataset. Experiments show that Robust-Wide can achieve an average BER of 0.96% on ControlNet-InstructPix2Pix and 9.34% on MagicBrush. Figure 4 shows some visual examples (more examples are provided in the supplementary material). For the instruction “swap sunflowers with roses", MagicBrush makes the edited image more different from the original image, leading to a relatively higher but still acceptable BER.
@@
We also tested the robustness of Robust-Wide against Inpainting [10] and DDIM Inversion[16] based on Stable Diffusion Models. We found that, even though Robust-Wide has never seen these image editing methods during training, it still demonstrates effective resistance as shown in Figure 4.
Continual Editing. A user may utilize InstructPix2Pix to perform multiple edits on a single image. An image edited by one user could also be spread to another for further editing, and this process can repeat several rounds. Hence, we need to ensure our watermarks can resist continual editing.
@@
As shown in Figure 5, we observe that the watermark embedded into the original image can be accurately extracted even after 3 editing rounds, demonstrating the watermark’s strong robustness against continual editing.